OpenStudy (lgbasallote):

400th question special: Proof that vampires do NOT exist..Calculus way (sorry twilight fans)

OpenStudy (lgbasallote):

Let v = number of vampires t = months Let's say the first vampire, Dracula, existed on the year 1600. Therefore, at t = 0, v = 1. Dracula sleeps for the whole month. Then, during the full moon, he turns someone into a vampire. Therefore, at t = 1, v = 2. Now, we use the formula \[\huge \ln (\frac{v}{v_o}) = k(t - t_o)\] where: \(v_o\) = initial number of vampires \(v\) = final number of vampires \(t\) = final time \(t_o\) = initial time \(k\) = constant so now, let us solve for the constant \[\huge \ln (\frac{2}{1}) = k(1 - 0)\] \[\huge \ln (2) = k\] \[\huge 0.6931 = k\] Now, let us calculate how many vampires there would be 10 years later or when t = 120 months. To solve for v, we use the same equation again. \[\huge \ln (\frac{v}{v_o}) = kt\] \[\huge \frac{v}{v_o} = e^{kt}\] \[\huge v = v_o e^{kt}\] Now, we substitute. \[\huge v = (1)[e^{(0.6931)(120)}]\] \[\huge v = 1.3292 \times 10^{36}\] Now, I doubt there were that huge amount of population in the world by the year 1610. Therefore, we can say that according to the formula for growth in Calculus, vampires do NOT exist!

OpenStudy (anonymous):

lol awesome

OpenStudy (anonymous):

Very amusing and I actually understood it :D

OpenStudy (anonymous):

VAMPIRES DEFY CALCULUS THEY DO EXIST!

OpenStudy (anonymous):

just like the talking chocolate m&m's

OpenStudy (lgbasallote):

wonder if "shiny" vampires also exist..hmm

OpenStudy (anonymous):

Im seriously lost.. hahah.. XD

OpenStudy (lgbasallote):

bottomline...vampires dont exist lol

OpenStudy (unklerhaukus):

dosent this model assume the vampires are imortal?

OpenStudy (lgbasallote):

this model assumes that everyone in the world by now are vampires

OpenStudy (unklerhaukus):

but surely some of the vampires are killed , maybe you need another term to give the population

OpenStudy (lgbasallote):

even if you cut half or even three fourths of that number...there's still too many

OpenStudy (unklerhaukus):

Professor Abraham Van Helsing

OpenStudy (lgbasallote):

hmm you're good -___- there's also Abraham Lincoln the Vampire Slayer

OpenStudy (unklerhaukus):

lolz

OpenStudy (anonymous):

This is a great way to start your day revising differential equations haha

OpenStudy (lgbasallote):

indeed it is haha

OpenStudy (anonymous):

How could you say, im a vampire. J/K Well it makes sense. the calculation is awesooome

OpenStudy (lgbasallote):

why thank you ;)

OpenStudy (wolfboy):

vlaud the impailer...the original dracula

OpenStudy (wolfboy):

But hey what about werewolves???

OpenStudy (anonymous):

However, vampires are not included in national censuses. They live in the shadows and only come out at night. Censuses are taken during the day time.

OpenStudy (lgbasallote):

actually Foster's Home for Imaginary Friends conduct censuses during nighttime

OpenStudy (anonymous):

But vampires aren't imaginary!

OpenStudy (anonymous):

All that you've proved is that many of the "people" I meet are actually vampires. In fact, I might be one and not know it.

OpenStudy (blockcolder):

Why has carrying capacity not been accounted for here?

OpenStudy (jamesj):

Argument by contradiction: Suppose vampires exist. Then because I have very tasty blood, I would be dead. But I am not. Therefore vampires don't exist. QED.

OpenStudy (lgbasallote):

that is a very good point..isnt that called a proof by contrapositive? or is it another thing?

OpenStudy (apoorvk):

Hmm, vampire's ain't living. So they're unreal. Imaginary. (existence in iota format ^.^). So as many of them may exist, but not necessary that they would occupy space in the world! There's is a world parallel to ours, called the Argand's World! (Count Argand was a great Vampire *sigh*. They lost him to the werewolves). So, your proof is-one-helluva-EPICFAIL!!!! (you have no idea, a vampy may just be standing right behind you. Reason why he isn't sucking your blood right now is that he's had his fill for tonight -.-) *MAY JUST BE* -.- "Concealed within the living world that we see, exists a dark.. unknown... unfathomable sphere of creatures that dwell on the blood of the real world... In this Argand's world, not a single drop of blood is wasted... *Welcome to DRACUVILLE!* "

OpenStudy (apoorvk):

*Their's

OpenStudy (apoorvk):

@lgbasallote the fundamental error due to which you got that huge no. is that you didn't account for Vampires being killed (by werewolves ofcourse) -.-

OpenStudy (anonymous):

GO TEAM EDWARD!!!!!

OpenStudy (apoorvk):

uhh, Edward was expelled from the group of vampires. He was too bad and disgraceful to our community. -.-

OpenStudy (apoorvk):

So don't even mention his name unless you want to be thrown to the werewolves like we did to him. -_-

OpenStudy (anonymous):

@lgbasallote that was a proof by contradiction, not by contrapositive. In a proof by contradiction (reductio ad absurdum) you begin by assuming the opposite of the statement that you hope to prove. If you hope to prove that sqrt(2) is irrational, you begin with the assumption that it's rational. If you hope to prove that objects fall down, then you begin with the assumption that they fall up. If you hope to prove that God exists, you begin with the assumption that God does not exist. Once you've made that assumption, you follow logic to some absurd, obviously false conclusion. As long as all of your logic is sound, you can conclude that your assumption was false. Since you assumed the opposite of what you wanted to prove, then you've proven your case by disproving the opposite of it. Proof by contrapositive is a little different. You still do a direct proof of some sort. You don't assume the opposite of your statement or anything, but instead begin with a normal assumption. However, what is different is that instead of proving the statement itself, you prove the contrapositive of the statement, which is logically equivalent. An example of this would be; I want to prove the statement "If a number is even, then it can be divided by 2." Instead, I could choose to prove the logically equivalent contrapositive, which is, "If a number cannot be divided by 2, then it is odd."